# Measuring Sample Importance in Data Pruning for Training LLMs from a Data Compression Perspective

### v1

@@ 2.1 Compression View to Data Pruning
We view $H(W,p)$ as related to the minimum description length required to compress words in $W$. Indeed, the optimal compression based on entropy, e.g, Huffman coding Huffman [1952] or arithmetic coding Rissanen [1976], uses approximately $\lceil\log_{2}1/p(w_{i}|w_{<i})\rceil$ bits to encode $w_{i}$ given previous words. Alternatively, $H(W,p)$ represents the amount of information per word in passage $W$. Consider a next-word predictor $q$, i.e., a language model. The negative log-likelihood of $W$ on model $q$, which is our main metric for sample importance, is given by
@@
From an information theoretic view, (1) represents the description length of words in $W$ using sub-optimal distribution $q$. Meanwhile, if we average $nH(W,q)$ over $W$, it is the log-loss used for training language models. Thus, training is equivalent to minimizing the description length of, i.e., compressing, the corpus data Deletang et al. [2023].
@@ 3.4 Downstream tasks
We finetune and evaluate each model in downstream tasks such as text classification and textual similarity tasks from glue Wang et al. [2018]. The top of Fig. 3 shows the text classification results of 125M models on CoLA dataset, which compare between the proposed method and random pruning. The results show that, the accuracy of our method is always above that of the random pruning. Next, we scaled up the target model to 345M, and the bottom of Fig. 3 shows that the proposed data pruning method achieves significant performance gains over random pruning. Notably, the proposed pruning outperforms no-pruning case with almost up to 50% pruning ratio, i.e., pruning helps improving the performance of language models.

### v2

@@ 2.1 Compression View to Data Pruning
We view $H(W,p)$ as related to the minimum description length required to compress words in $W$. Indeed, the optimal compression based on entropy, e.g, Huffman coding Huffman [1952] or arithmetic coding Rissanen [1976], uses approximately $\lceil\log_{2}1/p(w_{i}|w_{<i})\rceil$ bits to encode $w_{i}$ given previous words. Alternatively, $H(W,p)$ represents the amount of information per word in passage $W$. Consider a next-word predictor $q$, i.e., a language model. The negative log-likelihood of $W$ on model $q$ averaged over sequence length, which is our main metric for sample importance, is given by
@@
From an information theoretic view, (1) represents the description length of words in $W$ using sub-optimal distribution $q$. Meanwhile, if we average $nH(W,q)$ over $W$, it is the log-loss used for training language models. Thus, training a language model is equivalent to minimizing the description length of, i.e., compressing, corpus data, as observed in Deletang et al. [2023]. In practice, the samples in batches have a fixed sequence length when training Transformer-based causal language models such as GPT-3 Brown et al. [2020] or Llama-2 Touvron et al. [2023], where the sequence length $n$ is fixed to 2048 and 4096 respectively. For example, if the lengths of documents are larger than $n$, they are divided into multiple sequences (samples); if it is shorter than $n$, multiple documents are packed into a single sequence Raffel et al. [2020]. Thus, from the implementation point of view, we can regard sequence length $n$ as fixed, and (1) can be regarded as being proportional to the negative log-likelihood or log-loss.
@@ 3.4 Downstream tasks
We finetune and evaluate each model in downstream tasks such as text classification and textual similarity tasks from glue Wang et al. [2018]. The left of Fig. 3 shows the text classification results of 125M models on CoLA dataset, which compare between the proposed method and random pruning. The results show that, the accuracy of our method is always above that of the random pruning. Next, we scaled up the target model to 345M, and the right of Fig. 3 shows that the proposed data pruning method achieves significant performance gains over random pruning. Notably, the proposed pruning outperforms no-pruning case with almost up to 50% pruning ratio, i.e., pruning helps improving the performance of language models.